# Robust Matrix Completion for Discrete Rating-Scale Data

### v1

@@ 4 Empirical Applications
We consider two publicly available benchmark datasets in the context of recommender systems. The first is used to investigate the impact of attacks on empirical data (rather than stylized simulated data), while the second allows to compare two missing data mechanisms (MNAR and MCAR). In both applications, the observed ratings are split randomly into subsets of 80% for training the algorithms and 20% for testing prediction performance.
@@
Throughout this section, we also analyze the stability of RDMC and SI with respect to their stopping criteria, in order to assess whether reducing the computational burden comes at a cost of accuracy. We use two stopping criteria for each method, denoted by liberal and strict. For RDMC, the liberal criterion corresponds to stopping after a maximum of 10 iterations, whereas the strict criterion iterates until convergence (cf. our simulations from Section 3). For SI, on the other hand, we follow the strategy of Hastie et al., (2015) by setting a larger convergence threshold of $10^{-3}$ for the liberal criterion, while the strict criterion uses the threshold $10^{-4}$ (as in the simulations). The other parameters are kept the same as in Section 3, except that the number of replications in repeated holdout sampling for selecting the regularization parameter is increased from 5 to 10.
@@ 4.1 MovieLens 100K data: The impact of attacks
The famous MovieLens 100K dataset (Harper & Konstan,, 2015) is extensively analyzed in previous studies on recommender systems (e.g., Tang & Guan,, 2020; Mazumder et al.,, 2020; Wang et al.,, 2024). It consists of 100,000 movie ratings on a scale from 1 to 5, of $n=943$ users on 1682 movies. We restrict our analysis to the $p=939$ movies that are rated at least 20 times. The observed ratings follow a left-skewed distribution with a majority of medium to high ratings, as shown in Figure 5.
@@
First, Figure 6 displays the prediction performance of the different methods in terms of the MAE on the test set. RDMC (with any of the three loss functions) and SI clearly outperform median and mode imputation with an MAE of around 0.725, while SI-discretized presents an even smaller MAE around 0.69. Applying the liberal stopping criterion thereby hardly affects prediction performance of RDMC and SI compared to the strict criterion.
Second, we investigate the sensitivity of the methods to nuke attacks. To this end, we use the three attack schemes described in Section 3.2.3. We additionally vary the number of fake profiles, based on a proportion $\varepsilon\in\{0.10,0.15,0.20\}$ of the number of observed ratings in the target variable. Figure 7 displays the results for the MPS. All RDMC approaches outperform both SI approaches by a considerable margin, for all attack schemes. Moreover, RDMC provides effective protection against the attacks with an MPS close to 0 in almost all cases, irrespective of the loss function. In the presence of the love/hate attack, however, the strict stopping criterion is necessary for RDMC to achieve the best result. As expected, the MPS of the two SI approaches deteriorates as the size of the attack increases. Interestingly, SI-discretized falls behind SI when the attack size exceeds 15%.
Figure 6: Prediction performance for the MovieLens 100K data.
@@ 4.2 Yahoo! Music data: The impact of missingness mechanisms
The Yahoo! Music ratings data [^2] Different datasets are available through the Yahoo! Webscope data sharing program at https://webscope.sandbox.yahoo.com/catalog.php?datatype=r. We use the dataset titled “R3 - Yahoo! Music ratings for User Selected and Randomly Selected songs, version 1.0”, downloaded on September 9, 2024. contain two datasets of users rating songs on a scale of 1 to 5. In the first dataset (User Selected), the users choose the songs they want to rate, whereas in the second dataset (Randomly Selected), 10 songs from a pool of $p_{2}=1000$ are randomly assigned to them. Hence, the former constitutes an MNAR mechanism while the latter represents the MCAR mechanism. To ensure a fair comparison across datasets, we restrict our analysis to the $n=2361$ users who are present in both datasets and rated at least 20 songs in the first dataset. Furthermore, we limit the first dataset to the $p_{1}=998$ songs with at least 20 ratings. In both datasets, we have more than 95% of missing values (96% for MNAR and 99% for MCAR). Figure 8 displays the distributions of the observed ratings in the two datasets. The missing data mechanism clearly leads to different rating distributions. When the users can choose the songs they want to evaluate (MNAR, left plot), the distribution of the ratings is almost bimodal with 1-star ratings being the most frequent, followed by 5-star ratings. When the songs to be rated are randomly assigned (MCAR, right plot), the distribution is highly right-skewed with decreasing frequency for increasing ratings, and very few high ratings.
@@
Figure 9 shows the MAE on the test set for both missing data mechanisms. As expected, the MAE of all methods is lower in case of MCAR compared to MNAR. Overall, RDMC outperforms the SI approaches in both scenarios. For RDMC, the loss function and the number of iterations seem to have an effect only for the MNAR scenario, with a small advantage for the absolute loss function and perhaps a minor improvement with the liberal stopping criterion. For SI, the discretized variant indicates only a minimal improvement over the standard algorithm, but the stopping criterion has a major impact for the MNAR scenario: the strict stopping criterion is required to ensure performance similar to RDMC.
@@
5 Conclusions
Motivated by applications in recommender systems as well as surveys in the social and behavioral sciences, we present a method for robust matrix completion with discrete rating-scale data. The method thereby combines a robust loss function on the errors for the observed ratings with a discreteness constraint on the predictions and a low-rank constraint on an ancillary continuous matrix. Extensive simulations focused on missing not at random (MNAR) mechanisms demonstrate that the proposed method protects well against corrupted observations—such as attacks with fake user profiles on recommender systems or careless respondents in surveys—while paying only a small price in terms of performance in the absence of corrupted observations. Two empirical applications further highlight the practical relevance of the method.
For recommender systems applications, the choice among the three considered loss functions does not yield considerable differences in performance, with the pseudo-Huber loss being recommended overall due to in general requiring fewer iterations for convergence. Moreover, performing only a small number of iterations (rather than iterating until convergence) can reduce computational effort, in many cases without a detrimental effect on performance. Fur survey applications, the choice of loss function plays a more important role, with the pseudo-Huber loss being recommended if one expects relatively few careless respondents, whereas the bounded absolute loss is recommended if one expects a sizable number of careless respondents.
Computational details
All computations were performed in R (R Core Team,, 2024) with packages rdmc (Alfons & Archimbaud,, 2024) and softImpute (Hastie & Mazumder,, 2021). Replication files will be made publicly available upon acceptance of this manuscript.
@@ A.1 Simulations motivated by recommender systems
A.2 Simulations of survey data
@@
Figures A.4A.7 display the results for the simulated surveys for the settings where 40% and 60% of respondents abandon the survey. The results are qualitatively similar to the results discussed in the main text with 20% of respondents abandoning the survey.
@@
Figure A.4: Results for the simulated survey with $p=40$ variables and 40% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
@@
Figure A.5: Results for the simulated survey with $p=40$ variables and 60% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
@@
Figure A.6: Results for the simulated survey with $p=80$ variables and 40% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
@@
Figure A.7: Results for the simulated survey with $p=80$ variables and 60% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.

### v2

@@ 4 Empirical case studies
We consider two publicly available datasets in the context of recommender systems. The first in Section 4.1 is used to investigate the impact of attacks on empirical data (rather than stylized simulated data), while the second in Section 4.2 allows to compare two missing data mechanisms (MNAR and MCAR). In both case studies, the observed ratings are split randomly into subsets of 80% for training the algorithms and 20% for testing prediction performance.
@@
Since computational burden forms an important practical consideration across many recommender system applications, we also analyze the stability of RDMC and SI with respect to their stopping criteria in order to assess whether reducing the computational burden comes at a cost of accuracy. We use two stopping criteria for each method, denoted by liberal and strict. For RDMC, the liberal criterion corresponds to stopping after a maximum of 10 iterations, whereas the strict criterion iterates until convergence (cf. our simulations from Section 3). For SI, on the other hand, we follow the strategy of Hastie et al., (2015) by setting a larger convergence threshold of $10^{-3}$ for the liberal criterion, while the strict criterion uses the threshold $10^{-4}$ (as in the simulations). The other parameters are kept the same as in Section 3, except that the number of replications in repeated holdout sampling for selecting the regularization parameter is increased from 5 to 10.
@@ 4.1 MovieLens 100K data: The impact of attacks
The famous MovieLens 100K dataset (Harper & Konstan,, 2015) is extensively analyzed in previous studies on recommender systems (e.g., Mao et al.,, 2019; Tang & Guan,, 2020; Mazumder et al.,, 2020). It consists of 100,000 movie ratings on a scale from 1 to 5, of $n=943$ users on 1,682 movies. We restrict our analysis to the $p=939$ movies that are rated at least 20 times. The observed ratings follow a left-skewed distribution with a majority of medium to high ratings, as shown in Figure 3.
@@
First, Figure 4 (left) displays the prediction performance of the different methods in terms of the MAE on the test set. RDMC (with any of the three loss functions) and SI clearly outperform median and mode imputation with an MAE of around 0.725, while SI-discretized presents an even smaller MAE around 0.69. Applying the liberal stopping criterion thereby hardly affects prediction performance of RDMC and SI compared to the strict criterion. However, Figure 4 (right) shows that it yields a substantial reduction in computation time for RDMC—by more than a factor of three—resulting in similar computation time to SI.
Figure 4: Prediction performance (left panel) and computation time in seconds (right panel) for the MovieLens 100K data. Computation time is averaged over a grid of ten values for $\lambda$ and in case of RDMC also across the three considered loss functions.
@@
Second, we investigate the sensitivity of the methods to adversarial manipulation. To this end, we use the three nuke attack schemes described in Section 3.3. We additionally vary the number of fake profiles, based on a proportion $\varepsilon\in\{0.10,0.15,0.20\}$ of the number of observed ratings in the target variable. Figure 5 displays the results for the MPS. All RDMC approaches outperform both SI approaches by a considerable margin, for all attack schemes. Moreover, RDMC provides effective protection against the attacks with an MPS close to 0 in almost all cases, irrespective of the loss function. In the presence of the love/hate attack, however, the strict stopping criterion is necessary for RDMC to achieve the best result. As expected, the MPS of the two SI approaches deteriorates as the size of the attack increases. Interestingly, SI-discretized falls behind SI when the attack size exceeds 15%.
@@ 4.2 Yahoo! Music data: The impact of missingness mechanisms
The Yahoo! Music ratings data contain two datasets of users rating songs on a scale of 1 to 5. In the first dataset (User Selected), the users choose the songs they want to rate, whereas in the second dataset (Randomly Selected), 10 songs from a pool of $p_{2}=1,000$ are randomly assigned to them. Hence, the former constitutes an MNAR mechanism while the latter represents the MCAR mechanism. To ensure a fair comparison across datasets, we restrict our analysis to the $n=2,361$ users who are present in both datasets and rated at least 20 songs in the first dataset. Furthermore, we limit the first dataset to the $p_{1}=998$ songs with at least 20 ratings. In both datasets, we have more than 95% of missing values (96% for MNAR and 99% for MCAR). Figure 6 displays the distributions of the observed ratings in the two datasets. The missing data mechanism clearly leads to different rating distributions. When the users can choose the songs they want to evaluate (MNAR, left plot), the distribution of the ratings is almost bimodal with 1-star ratings being the most frequent, followed by 5-star ratings. When the songs to be rated are randomly assigned (MCAR, right plot), the distribution is highly right-skewed with decreasing frequency for increasing ratings, and very few high ratings.
@@
Figure 8: Computation time in seconds for the Yahoo datasets, averaged over a grid of ten values for $\lambda$ and in case of RDMC also across the three considered loss functions. The columns correspond to different missing data mechanisms.
Figure 7 shows the MAE on the test set for both missing data mechanisms. As expected, the MAE of all methods is lower in case of MCAR compared to MNAR. Overall, RDMC outperforms the SI approaches in both scenarios. For RDMC, the loss function and the number of iterations seem to have an effect only for the MNAR scenario, with a small advantage for the absolute loss function and perhaps a minor improvement with the liberal stopping criterion. For SI, the discretized variant indicates only a minimal improvement over the standard algorithm, but the stopping criterion has a major impact for the MNAR scenario: the strict stopping criterion is required to ensure performance similar to RDMC. Nevertheless, as illustrated in Figure 8, RDMC with the liberal stopping criterion remains competitive with SI in terms of computation time.
5 Conclusions and discussion
In a recent review paper, LeBlanc et al., (2024) conclude that too few statisticians are engaged in research on recommender systems despite their potential to make consequential contributions. But to fulfill this potential, it is vital that recommender systems are studied under more realistic scenarios. Our simulations and case studies may thereby provide a blueprint—or at least useful starting values—on how to design such studies.
Moreover, we present a novel method for robust matrix completion with discrete rating-scale data (RDMC), as well as an extension of the popular method Soft-Impute (SI) with a post-hoc discretization step (SI-discretized). RDMC thereby combines a robust loss function on the errors for the observed ratings with a discreteness constraint on the predictions and a low-rank constraint on an ancillary continuous matrix. Extensive simulations and two empirical case studies focused on attacks with fake user profiles and missing not at random (MNAR) mechanisms demonstrate that RDMC protects well against adversarial manipulation, while paying only a small price in terms of performance in the absence of corrupted observations compared to SI-discretized. The choice among the three considered loss functions does not yield considerable differences in performance, with the pseudo-Huber loss being recommended overall due to in general requiring fewer iterations for convergence. Moreover, performing only a small number of iterations (rather than iterating until convergence) can reduce computational effort, in many cases without a detrimental effect on performance. Regarding Soft-Impute, the discretization step results in an improvement over its standard variant in the absence of corrupted observations, in particular in settings with few rating categories (five or less), but it remains equally vulnerable to fake profiles.
Our method shows promise for broader applicability across diverse domains. For instance, empirical research in social and behavioral sciences relies heavily on rating-scale surveys, but listwise deletion remains the predominant method of handling missing data in some fields (e.g., Peng et al.,, 2023). Preliminary simulations suggest that RDMC represents a notable advancement in robust imputation of rating-scale data (see Appendix B), even in the presence of inattentive respondents or bot respondents—a critical issue in online surveys. Furthermore, our approach holds potential for contexts involving so-called strategic missing data situations (Zhang & Wang,, 2019)—such as financial reporting, college admissions, job applications, and marketing—where individuals may intentionally withhold information to achieve favorable outcomes. Finally, recent research has shown a notable rise in matrix completion methods for panel data (e.g. Athey et al.,, 2021; Bai & Ng,, 2021; Choi & Yuan,, 2024), providing a promising direction for future adaptation of our robust methodology.
Data availability
The MovieLens 100K data are publicly available from https://grouplens.org/datasets/movielens/100k/. The Yahoo! Music data are available through the Yahoo! Webscope data sharing program at https://webscope.sandbox.yahoo.com/catalog.php?datatype=r. Various datasets are available through this link; we use the dataset titled “R3 - Yahoo! Music ratings for User Selected and Randomly Selected songs, version 1.0”, downloaded on September 9, 2024.
@@ Appendix A Additional simulation results
Appendix B Rating-scale surveys
B.1 Introduction
Another relevant application of rating-scale data, although rarely considered in the literature on matrix completion, are surveys in the social and behavioral sciences. In such surveys, researchers measure each latent construct of interest (e.g., personality traits) by multiple rating-scale items (e.g., respondents may be asked how accurately certain adjectives describe their personality, with response categories ranging from 1 = “very accurate” to 5 = “very accurate”; Arias et al.,, 2020). Missing values are commonly occurring due to nonresponse in parts of the survey, with nonresponse often being missing not at random (MNAR). For instance, participants may abandon the survey yielding higher probability of missingness for later survey items, or they may not respond to items they find, e.g., confusing or inappropriate. In addition, careless responding (i.e., responses not based on item content due to inattention or misunderstanding; e.g., Ward & Meade,, 2023) or bot responding (e.g., Storozuk et al.,, 2020) are common occurrences in online surveys, and a prevalence as low as 5% can invalidate research findings (e.g., Credé,, 2010; Arias et al.,, 2020; Welz et al., 2024a, ; Welz et al., 2024b, ).
In some fields, removing participants with missing responses is still the most common method of handling missing responses in survey data (e.g., Peng et al.,, 2023). Yet this constitutes a loss of valuable information (the observed responses of the removed participants) and generally results in biased analyses (e.g., Little & Rubin,, 2019). Furthermore, the common presence of careless respondents or bot respondents requires imputation methods designed for discrete rating-scale data. We therefore investigate the use of the proposed matrix completion method RDMC in preliminary simulations motivated by surveys in the social and behavioral sciences.
B.2 Simulations
B.2.1 Data generation
We simulate rating-scale responses of $n=300$ participants to a survey on $q=10$ latent constructs, with each construct being measured by $r\in\{4,8\}$ items such that the total number of items is $p\in\{40,80\}$. For this purpose, we first simulate the underlying latent sentiments $\bm{Z}$ from a multivariate normal distribution $N(\bm{0},\bm{\Sigma})$, where $\bm{\Sigma}$ follows a block-Toeplitz structure with 1 on the diagonal and the remaining elements being given by $\rho_{kl}=0.6^{|k-l|+1}$ with $k=1,\dots,q$ and $l=1,\dots,q$ denoting the row and column indices of the blocks in $\bm{\Sigma}$.
In survey-based research, skewness towards one side of the rating scale is common, for instance, in organizational and management research (cf. Becker et al.,, 2019; Alfons et al.,, 2022). Hence, we generate items with skewed distributions by adding random mean shifts to the columns of $\bm{Z}$—with all items within the same construct receiving the same mean shift—before discretization into $K\in\{5,7,9\}$ answer categories interpreted as values $\{1,\dots,K\}$. Similar to the previous simulations, the mean shifts are drawn from the interval $[0,s_{\text{max}}]$, where $s_{\text{max}}$ is chosen to result in 40% of the responses being expected in the highest answer category.
The resulting continuous data matrix is then discretized using equispaced breakpoints $-K/2+1,-K/2+2,\dots,K/2-2,K/2-1$. To simulate reverse-keyed items, which are commonly used in practice to help detect response inattention, half the items of each construct are reversed by reassigning answer category 1 to $K$, 2 to $K-1$, and so on. Note that this implies that the data contains both left-skewed and right-skewed items. As is common in the behavioral sciences, we randomize the order of the items in the survey (with the same order being used for all participants), resulting in the final rating-scale data matrix $\bm{R}$.
B.2.2 Missing values and careless responding
While online data collection tools can be set up to require participants to respond to individual questions in order to progress in the survey, participants may abandon the survey altogether. We therefore inject missing values into the data matrix $\bm{R}$ by simulating respondents who stop responding at some point in the survey. We set the proportion of respondents who abandon the survey to $\gamma\in\{0.2,0.4,0.6\}$. For each of those respondents, we randomly select the item at which they stop responding, with all responses from this item onwards being replaced by missing values. We emphasize that here $\gamma$ does not indicate the proportion of missing cells in $\bm{R}$, but the number of rows that contain missing values.
To simulate careless respondents, a proportion $\varepsilon\in\{0,0.1,0.2\}$ of the rows in $\bm{R}$ are replaced with careless respondents who randomly select from the two extreme answer categories $\{1,l\}$. Although this is a common careless response style (e.g., Baumgartner & Steenkamp,, 2001; Alfons & Welz,, 2024), it may not be realistic that all careless respondents behave in this way. Yet we chose this careless response style because the high leverage of the responses should be particularly influential for matrix completion methods. Note that our simulation allows for the situation that a careless respondent abandons the survey such that a certain row may contain both careless responses and missing values.
B.2.3 Evaluation criteria
B.2.4 Results
@@
We focus on the results from Figures B.1 and  B.2 for the settings with 20% of respondents who abandon the survey, as the results for 40% and 60% are similar (see Figures B.3B.6).
Box plots for the MAE in the simulated survey with $p=40$ items are shown in Figure B.1. In general, differences between the methods are smaller than in the simulations for recommender systems. On the other hand, we now see more differences among RDMC with different loss functions. Most notably, RDMC with the pseudo-Huber function remains close to SI-discretized in the absence of careless responding, while RDMC with the truncated absolute loss yields the best results for 20% careless respondents. In addition, with a higher number of answer categories, robustness benefits of RDMC over SI become more apparent. For instance, with 10% of careless respondents, RDMC and SI-discretized perform similarly for five answer categories, whereas all three variants of RDMC outperform SI-discretized in case of seven or nine answer categories.
@@
Figure B.1: Results for the simulated survey with $p=40$ variables and 20% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Increasing the number of items to $p=80$, Figure B.2 contains the corresponding box plots of the MAE. We now observe bigger differences between the methods. For all settings regarding the number of answer categories, the MAE of RDMC remains in between that of the two SI approaches in case of no careless responding, while RDMC yields better performance in the presence of careless respondents. As previously, the pseudo-Huber loss is preferable in the absence of careless responding (with an MAE close to that of SI-discretized), while the truncated absolute loss is the most robust for larger prevalence of careless respondents.
@@
Figure B.2: Results for the simulated survey with $p=80$ variables and 20% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Finally, we included variants of RDMC with the three loss functions for which we perform at most 10 iterations. Here, the MAE is at best similar but in most cases slightly higher compared to iterating until convergence. For simplicity, we thus omit these variants from the figures.
In summary, the choice of loss function seems to play a more important role in a survey context than in our simulations of recommender systems (see Section 3). The pseudo-Huber loss is thereby recommended if one expects relatively few careless respondents, whereas the truncated absolute loss is recommended if one expects a sizable number of careless respondents.
@@
Figure B.3: Results for the simulated survey with $p=40$ variables and 40% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
@@
Figure B.4: Results for the simulated survey with $p=80$ variables and 40% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Figure B.5: Results for the simulated survey with $p=40$ variables and 60% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Figure B.6: Results for the simulated survey with $p=80$ variables and 60% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.